I need a brief explanation on how to solve system of linear equations using addition/elimination, I understand the other 2 but do not get this one please help!
3x+2y=-13
5x-1y=-13

Answer by ankor@dixie-net.com(22740)

3x + 2y = -13
5x - 1y = -13
We want to multiply one or both of the equations so that when we add them, one of the variables is eliminated
:
In this problem we just have to multiply the 2nd equation by 2, and add to the 1st
3x + 2y = -13
10x - 2y = -26
-------------------Adding eliminates y, find x
13x + 0 = -39
x = -39/13
x = -3
:
Find y using the 1st equation, replace x with -3
3(-3) + 2y = -13
-9 + 2y = -13
2y = -13 + 9
2y = -4
y = -4/2
y = -2
:
;
Check the solutions in the original 2nd equation
5x - 1y = -13
replace x with -3; replace y with -2
5(-3) - 1(-2) = -13
-15 + 2 = -13; confirms our solution of x=-3, y=-2